OpenStudy (anonymous):

(perform the indicated computations. write the answer in scientific notations) (4X106)−1

OpenStudy (anonymous):

the 6 and the -1 represent the exponent. I just copy and past :P

OpenStudy (anonymous):

I mean paste instead of the word past. lol

OpenStudy (anonymous):

1/4000000 = 2.5 × 10^-7

OpenStudy (anonymous):

OTL. sorry i mean. 4 * 10^-6

OpenStudy (anonymous):

How did get 2.5 part?

OpenStudy (anonymous):

no...actually i mean 2.5*10^-7. fail :(

OpenStudy (anonymous):

you get this --> 1/4000000 so that gives 2.5 * 10^(-7) because 1/4 = .25 .... now tack on 6 extra zeroes. to get .2500000000, then move to the right, so 2.5 * 10^-7

OpenStudy (anonymous):

oh I see now, thanks Ronncc.:)

OpenStudy (anonymous):

\[(2^{-1}x ^{-2}y ^{-1})^{-2}(2x ^{-3}y)^{-2}(16x ^{-3}y ^{3})^{0}/ (2x ^{-3}y ^{-5)^{2}}\]